A rectangular make-up case is 3 in. high & has a volume of 90 in.^3. The width is 1 in. less than the length. Find the length & width of the case.
Answer by Boreal(15235)

width=x
length=x+1
height=3
their product is 90 in^3
therefore 3x^2+3x=90
x^2+x=30, dividing by 3
x^2+x-30=0
(x+6)(x-5)=0
x=5 in , only sensible root
x+1=6 in
lwh=6*5*3=90 in^3